OpenStudy (anonymous):

Given: angle A is and angle C are supplementary; angle B and angle D are supplementary; angle A is congruent to angle B Prove: angle C is congruent to angle D

OpenStudy (anonymous):

If two angles are supplementary, they add up to 180 degrees. So, a + c = 180 b + d = 180 If angle a and angle b are congruent, we can substitute one for the other. Here, I will substitute a for b. a + c = 180 a + d = 180 We can subtract "a" from both sides of both equations. c = 180 - a d = 180 - a If angle c and angle d are equal to the same thing (180 - a), the angles themselves must have equal measures, and if they have equal measures, they are congruent. Hope I helped, and let me know if you need anything else or have any more questions! :)